OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the system. 3x – 7y = 11 2x + 5y = –4 Possible answers: 1 29 27 83

OpenStudy (amistre64):

how does your cramer rule look?

OpenStudy (amistre64):

3 11 2 -4 is the matrix for Dx right?

OpenStudy (amistre64):

or do we use the y parts for Dx? -7 11 5 -4 ???

OpenStudy (anonymous):

Dx

OpenStudy (jiteshmeghwal9):

first equal the value of 'x' in both of the equation as\[2(3x-7y)=(11)2\]\[3(2x+5y)=(-4)3\]\[\implies 6x-14y=22\]\[6x+15y=-12\]now change the signs in both of the equation

OpenStudy (jiteshmeghwal9):